# A scattering perspective on gravitational lensing

### v1

@@ Abstract
We note that this expression is not strictly valid in the Born regime, where the right-hand side vanishes. Additionally, for Coulomb potentials, one must regularize the divergences in the forward limit, $\theta=0$ Opt_Th_Coulomb_PhysRevLett.14.81; Review_Opt_Th_Coulomb.
@@
A natural next step is to include corrections within the post-Minkowskian (PM) framework Bertotti:1956pxu; Kerr:1959zlt; Bertotti:1960wuq; Westpfahl:1979gu; Portilla:1979xx; Westpfahl:1985tsl, which systematically accounts for weak-field relativistic corrections in scattering. In this setting, the lens can be systematically treated as a compact object with a well-defined gravitational potential given as a series expansion in $G$, allowing one to compute how scattering and lensing observables are modified beyond the Newtonian approximation. To provide this PM extension, we first note that, while the eikonal amplitude is obtained by resumming all the ladder and cross-ladder diagrams in its Feynman series expansion Abarbanel:1969ek; Levy:1969cr; tHooft:1987vrq, higher-order PM corrections to $\mathcal{M}^{\rm eik}$ arise from higher-order diagrams which are suppressed by powers of $GM|{\bm{q}}|$. For instance, the leading correction to Newton, i.e., the 2PM correction, is captured by the classical part of the one-loop triangle diagram.
In the following, we will analyze the particularly clean example of a point mass lens, described by the Schwarzschild solution. At 2PM order, the correction is given by the classical contribution of the Compton amplitude in the eikonal limit, $\omega b\gg 1$, which, after using Eq. (A scattering perspective on gravitational lensing), gives the potential
@@
This expression matches the Born series solution of a classical wave equation on a Schwarzschild background, as shown in Ref. Correia:2024jgr, and is in fact universal for non-spinning lenses. In this case, it is possible to show that the scattering amplitude exponentiates again, obtaining an expression with the same structure of Eq. (38), where the corresponding phase shift is computed as
@@
The second term in (A scattering perspective on gravitational lensing) therefore represents the first nonvanishing PM correction to the eikonal phase shift for a Schwarzschild-like lens. As expected, this contribution is suppressed by powers of $GM/|{\bm{b}}|$, and thus becomes relevant only when the lens lies sufficiently close to the emitting source.
@@
in terms of the Einstein ring measured from the lens position, $\tilde{\theta}_{\rm E}=(z_{\rm so}/z_{\rm lo})\theta_{\rm E}$. This parameter dictates the strength of the PM correction and, as easily appreciated, takes larger values for heavier lenses or for GW sources sufficiently close to the lens itself. To provide few representative values, it is possible to show that, for characteristic LISA binary systems located at distances $z_{\rm so}\simeq 100\,{\rm Mpc}$, intermediate-mass black hole lenses $M\sim 10^{3}M_{\odot}$, at distances $z_{\rm sl}\simeq{\rm pc}$, would give ${\cal C}|_{G^{2}}\approx\mathcal{O}(10^{-5})$. On the other hand, one can consider hierarchical triples where the source orbits, at a distance $z_{\rm sl}\simeq\mathcal{O}(10^{-3}\div 1)\,{\rm pc}$, a supermassive black hole lens with mass $M\sim\mathcal{O}(10^{6}\div 10^{8})M_{\odot}$ Antonini:2012ad; Berczik_2024. In this case, the PM coefficient can reach values as large as ${\cal C}|_{G^{2}}\approx\mathcal{O}(10^{-2})$.
Figure 2: Behavior of the modulus of the diffraction integral in terms of the dimensionless frequency $\nu$, including the PM corrections at order $\mathcal{O}(G^{2})$. Different colors show distinct values of the observer location $\varphi$ and the Einstein ring $\tilde{\theta}_{\rm E}$. The bottom panel is a zoom in of the top panel for $\nu\in(40,50)$.
@@
where $\varphi_{i}$ are the two images formed in a potential of a point-mass lens (without the PM-corrections), with the convention $\varphi_{1}>\varphi_{2}$. As an example, we discuss the central panel of Fig. 4. In this case, the observer position is set to $\varphi=1.5$, and using the lens equation (10), we find the position of the two images $\varphi_{1}=2.0$ and $\varphi_{2}=0.5$, yielding a period of $T_{\rm long}\approx 5.7/\tilde{\theta}_{\rm E}$. This is in perfect agreement with Fig. 4, where the red curve show a beating pattern with frequency $\approx 570$, while the blue curve has a period of $\approx 5700$. Importantly, from Fig. 4 and Eq. (48), we see that the value of $\tilde{\theta}_{\rm E}$ primarily fixes the frequency of the beating pattern (together with the image positions), rather than the amplitude of the percentage variation, which is mostly dictated by $\varphi$ (see the different panels of Fig. 4). Indeed, at $\varphi=1.5$, the correction reaches roughly $60\%$ (central panel of Fig. 4), independently of $\tilde{\theta}_{\rm E}$.
@@
Remarkably, the result of Eq. (A scattering perspective on gravitational lensing) has been recently generalized in Bini:2025ltr; Bini:2025bll (following inspiration from Ivanov:2025ozg) by including all orders in the PM expansion, with the corresponding phase shift in the eikonal limit given by
@@
Indeed, this result can be obtained as the WKB solution to the radial wave equation in a Schwarzschild lens potential, and it is equal to the on-shell action for a null geodesic. While a detailed study of the effects of the PM-resummed eikonal phase in gravitational lensing is left for future work, we expect these corrections to become relevant at large deflection angles. In particular, while the PM expansion of Eq. (A scattering perspective on gravitational lensing) cannot reproduce the multiple (relativistic) images that appear when the source is sufficiently compact, the fully resummed eikonal phase should, in principle, be able to do so. Indeed, computing the scattering angle $\chi=-\frac{\partial\delta}{\partial|{\bm{b}}|}$ shows that it diverges logarithmically as $\chi({\bm{b}})\sim\log(|{\bm{b}}|-b_{\rm cr})$ at the critical impact parameter $b_{\rm cr}=3\sqrt{3}\,GM$, corresponding to the shadow of the lens. In this regime, $\chi({\bm{b}})$ becomes a multi-valued function with a branch cut at the shadow radius, leading to an infinite number of stationary points of the lens equation—each associated with a distinct image corresponding to a null geodesic that orbits the lens multiple times. By contrast, the PM-expanded scattering angle $\chi\sim 1/|{\bm{b}}|+\mathcal{O}(1/|{\bm{b}}|^{2})$, derived from Eq. (A scattering perspective on gravitational lensing), yields only two such solutions (three when the 2PM correction is included), describing small-angle deflections. Thus, while the PM expansion captures the perturbative regime, only the fully resummed eikonal phase encodes the full image multiplicity. This strong field limit near the black hole shadow has been analyzed in Bozza:2001xd, and a partial wave treatment was carried out numerically in Nambu:2015aea, but Eq. (A scattering perspective on gravitational lensing) can provide a more thorough analysis. We leave a detailed investigation to future work.
Beyond eikonal corrections. The partial wave expansion in Eq. (35) is exact in the far distance limit, and it was our starting point for deriving the scattering amplitude in the eikonal limit, showing that this matches the diffraction integral. We can improve this discussion by including beyond-eikonal (BE) corrections in the large $\ell$ limit of Eq. (35), which correspond to finite-wavelength corrections of ${\cal O}(\ell^{-1})\sim{\cal O}(1/b\omega)$. These corrections can arise from two sources: the large $\ell$ expansions of the phase shift $\delta^{\rm eik}$ and of the Legendre polynomials. The former can be extracted for a relativistic point-mass lens from the far-zone scattering phase shifts computed in black hole perturbation theory. For instance the phase shift for a spin-$s$ wave is given to 2PM order by Casals:2015nja; Bautista:2023sdf; Ivanov:2024sds; Caron-Huot:2025tlq
@@
The first line agrees with the eikonal phase shift in Eq. (A scattering perspective on gravitational lensing) including its 2PM correction, and is independent of spin, in accordance with the equivalence principle [^10] At the level of the partial-wave expansion, a further distinction between scalar and gravitational waves arises from the use of spin-weighted rather than scalar spherical harmonics, but this difference disappears in the eikonal limit for lenses which preserve helicity.. The second line corresponds to sub-eikonal ${\cal O}(1/\ell^{2})$ corrections, which depend on the spin of the wave. The correction at ${\cal O}(G)$ is purely Newtonian, whereas the second order one is a relativistic correction.
@@
The impact of the beyond-eikonal corrections on the diffraction integral is shown in Fig. 5, where we plot, for a fixed impact parameter $\varphi=1.5$, the percentage corrections to the amplitude of the amplification factor when including either the PM correction (red curve, ${\cal C}|_{\rm b\text{-}eik}=0$) or the beyond-eikonal correction (blue curve, ${\cal C}|_{G^{2}}=0$) alone. In contrast to the PM corrections, the trend of the beyond-eikonal term is to decrease for larger values of $\nu$, as expected from Eq. (54). Therefore, one expects the dominant source of corrections at large $\nu$ to arise from the PM terms, while at low $\nu$ the beyond-eikonal corrections should, in principle, dominate. In practice, in our case, the beyond-eikonal contributions are always negligible compared to the PM corrections. This suppression is a consequence of the specific gravitational potential adopted here: for a Newtonian potential, the first beyond-eikonal correction vanishes Bjerrum-Bohr:2019kec, [^11] Let us mention that relaxing the assumption of spherical symmetry for the lens can in principle generate a nonvanishing leading beyond-eikonal correction, although such contributions are expected to be PM suppressed. and the leading contribution scales as $\propto 1/\ell^{2}$, as shown in Eq. (51). Consequently, ${\cal C}_{G^{2}}\propto\tilde{\theta}_{\rm E}$, whereas ${\cal C}|_{\rm b\text{-}eik}\propto\tilde{\theta}_{\rm E}^{2}$, leading to the additional suppression seen in Fig. 5.
An additional source of finite-wavelength corrections arises from the large-$\ell$ approximation of the partial waves. We show this by considering the partial-wave expansion of the outgoing wave in Eq. (A scattering perspective on gravitational lensing) as, [^12] To get this result, one has to consider the large distance and small angle limits, obtained by approximating the Hankel functions as $h^{(1)}_{\ell}(\omega r)h^{(1)}_{\ell}(\omega r_{\rm s})\approx(-1)^{\ell+1}\frac{e^{i\omega r}e^{\frac{i(\ell+1/2)^{2}}{2\omega r}}}{\omega r}\frac{e^{i\omega z_{\rm sl}}e^{\frac{i(\ell+1/2)^{2}}{2\omega z_{\rm sl}}}}{\omega z_{\rm sl}}$ $\approx(-1)^{\ell+1}\frac{z_{\rm so}}{\omega z_{\rm sl}z_{\rm lo}}e^{\frac{i\omega{\bm{b}}_{\rm o}^{2}}{2}\frac{z_{\rm sl}}{z_{\rm so}z_{\rm lo}}}e^{\frac{i(\ell+1/2)^{2}}{2\omega}\frac{z_{\rm so}}{z_{\rm sl}z_{\rm lo}}}\frac{e^{i\omega|{\bm{r}}-{\bm{r}}_{s}|}}{\omega|{\bm{r}}-{\bm{r}}_{s}|}\,$, where we have included the Fresnel factor as discussed below Eq. (34).
@@
Using Hilb’s asymptotic formula, the Legendre polynomial at large angular momenta can be written as Amado:1985zz
@@
which is valid for small scattering angles, $\theta\ll 1/\ell$, with an error of order ${\cal O}(\ell^{-2})$, extending the approximation used to derive Eq. (37) to the leading sub-eikonal correction in $1/\ell$. Looking at Eq. (A scattering perspective on gravitational lensing), it is clear that the amplification factor is given by the parenthesis in the second line (see Eq. (2)). We then approximate the sum over $\ell$ as an integral over the impact parameter, improving the $P_{\ell}(\cos\theta)$ expansion as in Eq. (56), finding
@@
The percentage correction described by Eq. (A scattering perspective on gravitational lensing) is plotted in Fig. 6 as a function of frequency for two choices of observer positions $\zeta=\theta/\tilde{\theta}_{\rm F}=\{0.05,0.1\}$, and two choices of Fresnel scale $\tilde{\theta}_{\rm F}=\{0.001,0.01\}$. The plot shows that the $\mathcal{O}(1/\ell^{2})$ correction is expected to give a small contribution to the full diffraction integral.
@@
We have shown explicitly that the diffraction integral is equivalent to the eikonal limit of the scattering amplitude for waves propagating in a weak gravitational potential, upon including finite-distance effects encoded in the Fresnel factor. This observation explains why the diffraction‐integral formalism smoothly interpolates between geometric and wave‐optics regimes and provides a clear organizing principle for systematic extensions. In particular, we have demonstrated how Born-like expansions, partial-wave techniques, and PM (i.e., relativistic) corrections furnish controlled generalizations of the standard treatment. Within this unified framework, PM corrections are especially relevant in the geometric-optics regime, where the wavelength is much smaller than the lens scale and higher-order terms refine the classical deflection, image structure, and time-delay predictions. Conversely, beyond-eikonal corrections could become important in the wave-optics regime, where diffraction and interference dominate, and subleading corrections to the eikonal phase capture finite-wavelength and off-stationary-path effects. While in the case of a Newtonian point particle, these corrections are largely suppressed, we note that a leading beyond-eikonal correction can also be induced in theories beyond General Relativity with screening mechanisms, in which the potential in the screened regions is no longer the usual Newtonian one Joyce:2014kja; Brax:2020ujo; Carrillo-Gonzalez:2021mqj. Together, these extensions provide a systematic and improvable route to make accurate lensing predictions across all frequency regimes, yielding a compact and versatile framework that links GW lensing to well-developed scattering-amplitude techniques and enables controlled analytic modeling of lensing signatures in forthcoming GW observations.
@@
Looking forward, the scattering-based formulation naturally interfaces with effective field theories, offering a controlled avenue for extending current models. Indeed, an effective theory description of the process allows one to go beyond the point-particle approximation and incorporate the spin Kubota:2023dlz; Kubota:2024zkv and finite-size effects of realistic lenses, as well as to account for non-elastic phenomena such as wave absorption and polarization, which become particularly relevant when considering images close to black hole shadows at large frequencies (for a recent work along this direction see Chan:2025wgz). At the same time, the lensing perspective may allow for novel approaches to scattering problems: saddle-point structures, multiple-image formation, and multi-lens configurations provide concrete, physically realized examples of semiclassical propagation in nontrivial potentials, potentially informing broader developments in high-energy and gravitational scattering theory. We plan to investigate these avenues in future work.
Acknowledgments. A.G. acknowledges Job Feldbrugge and Sunao Sugiyama for generously providing the code that enabled the numerical calculations presented in this work and for helpful discussions. M.C.G is supported by the Imperial College Research Fellowship. V.D.L. is supported by NSF Grants No. AST-2307146, No. PHY-2513337, No. PHY-090003, and No. PHY-20043, by NASA Grant No. 21-ATP21-0010, by John Templeton Foundation Grant No. 62840, by the Simons Foundation [MPS-SIP-00001698, E.B.], by the Simons Foundation International [SFI-MPS-BH-00012593-02], and by Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (https://www.arch.jhu.edu/), which is supported by the NSF Grant No. OAC-1920103. A.G. is supported by funds provided by the Center for Particle Cosmology at the University of Pennsylvania. The work of M.T. is supported in part by the DOE (HEP) Award DE-SC0013528. JPM thanks Donato Bini for discussions and comments.

### v2

@@ Abstract
We note that this expression is not strictly valid in the Born regime, where the right-hand side vanishes. Additionally, for Coulomb potentials, one must regularize the divergences in the forward limit, $\theta=0$ Holdeman and Thaler (1965); Canto et al. (2020).
@@
A natural next step is to include corrections within the post-Minkowskian (PM) framework Bertotti (1956); Kerr (1959); Bertotti and Plebanski (1960); Westpfahl and Goller (1979); Portilla (1979); Westpfahl (1985), which systematically accounts for weak-field relativistic corrections in scattering. In this setting, the lens can be systematically treated as a compact object with a well-defined gravitational potential given as a series expansion in $G$, allowing one to compute how scattering and lensing observables are modified beyond the Newtonian approximation. To provide this PM extension, we first note that, while the eikonal amplitude is obtained by resumming all the ladder and cross-ladder diagrams in its Feynman series expansion Abarbanel and Itzykson (1969); Levy and Sucher (1969); ’t Hooft (1987), higher-order PM corrections to $\mathcal{M}^{\rm eik}$ arise from higher-order diagrams which are suppressed by powers of $GM|{\bm{q}}|$. For instance, the leading correction to Newton, i.e., the 2PM correction, is captured by the classical part of the one-loop triangle diagram.
In the following, we will analyze the particularly clean example of a point mass lens, described by the Schwarzschild solution. At 2PM order, the correction is given by the classical contribution of the Compton amplitude in the eikonal limit, $\omega b\gg 1$, which, after using Eq. (Scattering perspective on gravitational lensing), gives the potential
@@
This expression matches the Born series solution of a classical wave equation on a Schwarzschild background, as shown in Ref. Correia and Isabella (2025), and is in fact universal for non-spinning lenses. In this case, it is possible to show that the scattering amplitude exponentiates again, obtaining an expression with the same structure of Eq. (38), where the corresponding phase shift is computed as
@@
The second term in (Scattering perspective on gravitational lensing) therefore represents the first nonvanishing PM correction to the eikonal phase shift for a Schwarzschild-like lens. As expected, this contribution is suppressed by powers of $GM/|{\bm{b}}|$, and thus becomes relevant only when the lens lies sufficiently close to the emitting source.
@@
in terms of the Einstein ring measured from the lens position, $\tilde{\theta}_{\rm E}=(z_{\rm so}/z_{\rm lo})\theta_{\rm E}$. This parameter dictates the strength of the PM correction and, as easily appreciated, takes larger values for heavier lenses or for GW sources sufficiently close to the lens itself. To provide few representative values, it is possible to show that, for characteristic LISA binary systems located at distances $z_{\rm so}\simeq 100\,{\rm Mpc}$, intermediate-mass black hole lenses $M\sim 10^{3}M_{\odot}$, at distances $z_{\rm sl}\simeq{\rm pc}$, would give ${\cal C}|_{G^{2}}\approx\mathcal{O}(10^{-5})$. On the other hand, one can consider hierarchical triples where the source orbits, at a distance $z_{\rm sl}\simeq\mathcal{O}(10^{-3}\div 1)\,{\rm pc}$, a supermassive black hole lens with mass $M\sim\mathcal{O}(10^{6}\div 10^{8})M_{\odot}$ Antonini and Perets (2012); Berczik et al. (2024). In this case, the PM coefficient can reach values as large as ${\cal C}|_{G^{2}}\approx\mathcal{O}(10^{-2})$.
Figure 2: Behavior of the modulus of the diffraction integral in terms of the dimensionless frequency $\nu$, including the PM corrections at order $\mathcal{O}(G^{2})$. Different colors show distinct values of the observer location $\varphi$ and the Einstein ring $\tilde{\theta}_{\rm E}$. The bottom panel is a close-up of the top panel for $\nu\in(40,50)$.
@@
where $\varphi_{i}$ are the two images formed in a potential of a point-mass lens (without the PM-corrections), with the convention $\varphi_{1}>\varphi_{2}$. [^10] Notice that, when $\varphi_{1}\approx\varphi_{2}$, the period of the modulation becomes very large, effectively suppressing the effects of the modification. This is the case when $\varphi=0$, namely when source-lens-observer are all aligned. As an example, we discuss the central panel of Fig. 4. In this case, the observer position is set to $\varphi=1.5$, and using the lens equation (10), we find the position of the two images $\varphi_{1}=2.0$ and $\varphi_{2}=0.5$, yielding a period of $T_{\rm long}\approx 5.7/\tilde{\theta}_{\rm E}$. This is in perfect agreement with Fig. 4, where the red curve shows a beating pattern with frequency $\approx 570$, while the blue curve has a period of $\approx 5700$. Importantly, from Fig. 4 and Eq. (48), we see that the value of $\tilde{\theta}_{\rm E}$ primarily fixes the frequency of the beating pattern (together with the image positions), rather than the amplitude of the percentage variation, which is mostly dictated by $\varphi$ (see the different panels of Fig. 4). Indeed, at $\varphi=1.5$, the correction reaches roughly $60\%$ (central panel of Fig. 4), independently of $\tilde{\theta}_{\rm E}$.
@@
Remarkably, the result of Eq. (Scattering perspective on gravitational lensing) has been recently generalized in Bini and Di Russo (2025); Bini et al. (2025) (following inspiration from Ivanov et al. (2025)) by including all orders in the PM expansion, with the corresponding phase shift in the eikonal limit given by
@@
Indeed, this result can be obtained as the WKB solution to the radial wave equation in a Schwarzschild lens potential, and it is equal to the on-shell action for a null geodesic. While a detailed study of the effects of the PM-resummed eikonal phase in gravitational lensing is left for future work, we expect these corrections to become relevant at large deflection angles. In particular, while the PM expansion of Eq. (Scattering perspective on gravitational lensing) cannot reproduce the multiple (relativistic) images that appear when the source is sufficiently compact, the fully resummed eikonal phase should, in principle, be able to do so. Indeed, computing the scattering angle $\chi=-\frac{\partial\delta}{\partial|{\bm{b}}|}$ shows that it diverges logarithmically as $\chi({\bm{b}})\sim\log(|{\bm{b}}|-b_{\rm cr})$ at the critical impact parameter $b_{\rm cr}=3\sqrt{3}\,GM$, corresponding to the shadow of the lens. In this regime, $\chi({\bm{b}})$ becomes a multi-valued function with a branch cut at the shadow radius, leading to an infinite number of stationary points of the lens equation—each associated with a distinct image corresponding to a null geodesic that orbits the lens multiple times. By contrast, the PM-expanded scattering angle $\chi\sim 1/|{\bm{b}}|+\mathcal{O}(1/|{\bm{b}}|^{2})$, derived from Eq. (Scattering perspective on gravitational lensing), yields only two such solutions (three when the 2PM correction is included), describing small-angle deflections. Thus, while the PM expansion captures the perturbative regime, only the fully resummed eikonal phase encodes the full image multiplicity. This strong field limit near the black hole shadow has been analyzed in Bozza et al. (2001), and a partial wave treatment was carried out numerically in Nambu and Noda (2016), but Eq. (Scattering perspective on gravitational lensing) can provide a more thorough analysis. We leave a detailed investigation to future work.
Beyond eikonal corrections. The partial wave expansion in Eq. (35) is exact in the far distance limit, and it was our starting point for deriving the scattering amplitude in the eikonal limit, showing that this matches the diffraction integral. We can improve this discussion by including beyond-eikonal (BE) corrections in the large $\ell$ limit of Eq. (35), which correspond to finite-wavelength corrections of ${\cal O}(\ell^{-1})\sim{\cal O}(1/b\omega)$. These corrections can arise from two sources: the large $\ell$ expansions of the phase shift $\delta^{\rm eik}$ and of the Legendre polynomials. The former can be extracted for a relativistic point-mass lens from the far-zone scattering phase shifts computed in black hole perturbation theory. For instance the phase shift for a spin-$s$ wave is given to 2PM order by Casals and Ottewill (2015); Bautista et al. (2024); Ivanov et al. (2024); Caron-Huot et al. (2025)
@@
The first line agrees with the eikonal phase shift in Eq. (Scattering perspective on gravitational lensing) including its 2PM correction, and is independent of spin, in accordance with the equivalence principle [^11] At the level of the partial-wave expansion, a further distinction between scalar and gravitational waves arises from the use of spin-weighted rather than scalar spherical harmonics, but this difference disappears in the eikonal limit for lenses which preserve helicity.. The second line corresponds to sub-eikonal ${\cal O}(1/\ell^{2})$ corrections, which depend on the spin of the wave. The correction at ${\cal O}(G)$ is purely Newtonian, whereas the second order one is a relativistic correction.
@@
The impact of the beyond-eikonal corrections on the diffraction integral is shown in Fig. 5, where we plot, for a fixed impact parameter $\varphi=1.5$, the percentage corrections to the amplitude of the amplification factor when including either the PM correction (red curve, ${\cal C}|_{\rm b\text{-}eik}=0$) or the beyond-eikonal correction (blue curve, ${\cal C}|_{G^{2}}=0$) alone. In contrast to the PM corrections, the trend of the beyond-eikonal term is to decrease for larger values of $\nu$, as expected from Eq. (54). Therefore, one expects the dominant source of corrections at large $\nu$ to arise from the PM terms, while at low $\nu$ the beyond-eikonal corrections should, in principle, dominate. In practice, in our case, the beyond-eikonal contributions are always negligible compared to the PM corrections. This suppression is a consequence of the specific gravitational potential adopted here: for a Newtonian potential, the first beyond-eikonal correction vanishes Bjerrum-Bohr et al. (2020), [^12] Let us mention that relaxing the assumption of spherical symmetry for the lens can in principle generate a nonvanishing leading beyond-eikonal correction, although such contributions are expected to be PM suppressed. and the leading contribution scales as $\propto 1/\ell^{2}$, as shown in Eq. (51). Consequently, ${\cal C}_{G^{2}}\propto\tilde{\theta}_{\rm E}$, whereas ${\cal C}|_{\rm b\text{-}eik}\propto\tilde{\theta}_{\rm E}^{2}$, leading to the additional suppression seen in Fig. 5.
An additional source of finite-wavelength corrections arises from the large-$\ell$ approximation of the partial waves. We show this by considering the partial-wave expansion of the outgoing wave in Eq. (Scattering perspective on gravitational lensing) as, [^13] To get this result, one has to consider the large distance and small angle limits, obtained by approximating the Hankel functions as $h^{(1)}_{\ell}(\omega r)h^{(1)}_{\ell}(\omega r_{\rm s})\approx(-1)^{\ell+1}\frac{e^{i\omega r}e^{\frac{i(\ell+1/2)^{2}}{2\omega r}}}{\omega r}\frac{e^{i\omega z_{\rm sl}}e^{\frac{i(\ell+1/2)^{2}}{2\omega z_{\rm sl}}}}{\omega z_{\rm sl}}$ $\approx(-1)^{\ell+1}\frac{z_{\rm so}}{\omega z_{\rm sl}z_{\rm lo}}e^{\frac{i\omega{\bm{b}}_{\rm o}^{2}}{2}\frac{z_{\rm sl}}{z_{\rm so}z_{\rm lo}}}e^{\frac{i(\ell+1/2)^{2}}{2\omega}\frac{z_{\rm so}}{z_{\rm sl}z_{\rm lo}}}\frac{e^{i\omega|{\bm{r}}-{\bm{r}}_{s}|}}{\omega|{\bm{r}}-{\bm{r}}_{s}|}\,$, where we have included the Fresnel factor as discussed below Eq. (34).
@@
Using Hilb’s asymptotic formula, the Legendre polynomial at large angular momenta can be written as Amado et al. (1985)
@@
which is valid for small scattering angles, $\theta\ll 1/\ell$, with an error of order ${\cal O}(\ell^{-2})$, extending the approximation used to derive Eq. (37) to the leading sub-eikonal correction in $1/\ell$. Looking at Eq. (Scattering perspective on gravitational lensing), it is clear that the amplification factor is given by the parenthesis in the second line [see Eq. (2)]. We then approximate the sum over $\ell$ as an integral over the impact parameter, improving the $P_{\ell}(\cos\theta)$ expansion as in Eq. (56), finding
@@
The percentage correction described by Eq. (Scattering perspective on gravitational lensing) is plotted in Fig. 6 as a function of frequency for two choices of observer positions $\zeta=\theta/\tilde{\theta}_{\rm F}=\{0.05,0.1\}$, and two choices of Fresnel scale $\tilde{\theta}_{\rm F}=\{0.001,0.01\}$. The plot shows that the $\mathcal{O}(1/\ell^{2})$ correction is expected to give a small contribution to the full diffraction integral.
@@
We have shown explicitly that the diffraction integral is equivalent to the eikonal limit of the scattering amplitude for waves propagating in a weak gravitational potential, upon including finite-distance effects encoded in the Fresnel factor. This observation explains why the diffraction‐integral formalism smoothly interpolates between geometric and wave‐optics regimes and provides a clear organizing principle for systematic extensions. In particular, we have demonstrated how Born-like expansions, partial-wave techniques, and PM (i.e., relativistic) corrections furnish controlled generalizations of the standard treatment. Within this unified framework, PM corrections are especially relevant in the geometric-optics regime, where the wavelength is much smaller than the lens scale and higher-order terms refine the classical deflection, image structure, and time-delay predictions. Conversely, beyond-eikonal corrections could become important in the wave-optics regime, where diffraction and interference dominate, and subleading corrections to the eikonal phase capture finite-wavelength and off-stationary-path effects. While in the case of a Newtonian point particle, these corrections are largely suppressed, we note that a leading beyond-eikonal correction can also be induced in theories beyond general relativity with screening mechanisms, in which the potential in the screened regions is no longer the usual Newtonian one Joyce et al. (2015); Brax et al. (2020); Carrillo-González et al. (2021). Together, these extensions provide a systematic and improvable route to make accurate lensing predictions across all frequency regimes, yielding a compact and versatile framework that links GW lensing to well-developed scattering-amplitude techniques and enables controlled analytic modeling of lensing signatures in forthcoming GW observations.
@@
Looking forward, the scattering-based formulation naturally interfaces with effective field theories, offering a controlled avenue for extending current models. Indeed, an effective theory description of the process allows one to go beyond the point-particle approximation and incorporate the spin Kubota et al. (2024b, a) and finite-size effects of realistic lenses, as well as to account for non-elastic phenomena such as wave absorption and polarization, which become particularly relevant when considering images close to black hole shadows at large frequencies (for a recent work along this direction see Chan et al. (2025)). At the same time, the lensing perspective may allow for novel approaches to scattering problems: saddle-point structures, multiple-image formation, and multi-lens configurations provide concrete, physically realized examples of semiclassical propagation in nontrivial potentials, potentially informing broader developments in high-energy and gravitational scattering theory. We plan to investigate these avenues in future work.
Acknowledgments. A.G. acknowledges Job Feldbrugge and Sunao Sugiyama for generously providing the code that enabled the numerical calculations presented in this work and for helpful discussions. J.P.M. thanks Donato Bini for discussions and comments. We also thank Jose Diego for useful comments on the draft. M.C.G is supported by the Imperial College Research Fellowship. V.D.L. is supported by NSF Grants No. AST-2307146, No. PHY-2513337, No. PHY-090003, and No. PHY-20043, by NASA Grant No. 21-ATP21-0010, by John Templeton Foundation Grant No. 62840, by the Simons Foundation [MPS-SIP-00001698, Emanuele Berti], by the Simons Foundation International [SFI-MPS-BH-00012593-02], and by Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (https://www.arch.jhu.edu/), which is supported by the NSF Grant No. OAC-1920103. A.G. is supported by funds provided by the Center for Particle Cosmology at the University of Pennsylvania. The work of M.T. is supported in part by the DOE (HEP) Award DE-SC0013528.